OpenStudy (anonymous):

Solve : lim[ln(x)-ln^2(x)]^3 , x->e

OpenStudy (ash2326):

We have \[\lim _{x \to e} (\ln x- (\ln x)^2)^3\] we know ln e=1 so \[ (1- (1)^2)^3=0\] \[\lim _{x \to e} (\ln x- (\ln x)^2)^3=0\]

OpenStudy (anonymous):

Thank you!

OpenStudy (ash2326):

Welcome :D :D